OpenStudy (anonymous):

Maggie uses a credit card with a 17.4% APR compounded monthly to pay for car repairs totaling $843.12. She can pay $330 per month on the card. What will the total cost of this purchase be?

OpenStudy (tkhunny):

i = 0.174 j = I/12 = 0.0145 843.12*(1+j) - 330 = ??? ???*(1+j) - 330 = ?? ??*(1+j) - 330 = ?

OpenStudy (anonymous):

can you explain it? i didnt get an answer